Circular No. 9002

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 X2 (LINEAR) G. Hug, Scranton, KS, U.S.A., reports his recovery of comet P/2004

CB (cf. IAUC 8314) as an essentially stellar object on his CCD images obtained with a 0.56-m reflector. The indicated correction to the prediction on MPC 56803 is ΔT = +0.16 day.

2008 UT  α_{2000}  δ_{2000}  Mag.
Dec. 4.38549  9^{h}38^{m}28.^{s}71  +34^{o}42′44′′.7  19.7
4.38954  9 38 29.10  +34 42 47.1  19.7
4.39441  9 38 29.57  +34 42 51.1  19.3
4.39927  9 38 30.05  +34 42 53.8  19.5
4.40333  9 38 30.48  +34 42 56.9  19.5
5.29700  9 40 01.34  +34 53 43.8  19.5
5.32912  9 40 04.52  +34 54 07.5  19.7
5.33805  9 40 05.39  +34 54 14.1  19.3
5.35773  9 40 07.34  +34 54 28.7  19.6
5.36422  9 40 07.98  +34 54 33.3  19.6

The following orbital elements by B. G. Marsden, Smithsonian Astrophysical Observatory, are from 338 observations, 2004–2008 (mean residual 0′′.4):

Epoch = 2004 Mar. 16.0 TT

T = 2004 Apr. 2.1633 TT  ω = 149.^{o}6559
e = 0.689403  Ω = 66.4876 }2000.0
q = 0.912180 AU  i = 19.1473

a = 2.936863 AU n^{o} = 0.1958297 P = 5.03 years Epoch = 2009 Mar. 30.0 TT

T = 2009 Apr. 15.9744 TT  ω = 149.^{o}7295
e = 0.688978  Ω = 66.4497 }2000.0
q = 0.913706 AU  i = 19.1479

a = 2.937753 AU n^{o} = 0.1957407 P = 5.04 years

COMET 29P/SCHWASSMANN-WACHMANN Visual total-magnitude estimates (cf. IAUC 8991): Oct. 11.10 UT,

10.3 (M. Meyer, Frauenstein, Germany, 15-cm refractor); 26.11, 10.4 (J. J. Gonzalez, Leon, Spain, 20-cm reflector); Nov. 6.03, 10.3 (Gonzalez); 10.16, 11.0 (M. L. Paradowski, Dabrowa, Poland, 25-cm reflector); 23.83, 11.0 (S. Yoshida, Gunma, Japan, 40-cm reflector).

2008 December 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green